# First Hitting Times on a Quantum Computer: Topological Effects and Dark States

### 24 Feb 2024 v1

@@ Abstract
We investigate a quantum walk on a ring represented by a directed triangle graph with complex edge weights and monitored at a constant rate until the quantum walker is detected. To this end, the first hitting time statistics is recorded using unitary dynamics interspersed stroboscopically by measurements, which is implemented on IBM quantum computers with a midcircuit readout option. Unlike classical hitting times, the statistical aspect of the problem depends on the way we construct the measured path, an effect that we quantify experimentally. First, we experimentally verify the theoretical prediction that the mean return time to a target state is quantized, with abrupt discontinuities found for specific sampling times and other control parameters, which has a well-known topological interpretation. Second, depending on the initial state, system parameters, and measurement protocol, the detection probability can be less than one or even zero, which is related to dark-state physics. Both, return-time quantization and the appearance of the dark states are related to degeneracies in the eigenvalues of the unitary time evolution operator. We conclude that, for the IBM quantum computer under study, the first hitting times of monitored quantum walks are resilient to noise. Yet, finite resolution effects lead to new topological, chirality, and broadening effects, which disappear in the asymptotic theory with an infinite number of measurements. Our results point the way for the development of novel quantum walk algorithms that exploit measurement-induced effects on quantum computers.
@@ I Introduction
The option of midcircuit readout of qubits on state-of-the-art quantum computers [1] opens opportunities to test the dynamics and statistics of monitored quantum dynamics. The repeated monitoring at predetermined times yields a string of measurement outputs. These can be viewed as a stochastic trajectory varying in time which depends on many factors, including the initial state, the dynamics of the measurement-free process, or the chosen time intervals between measurements. Given a stochastic trajectory, the first quantum hitting time, defined as the time taken for a monitored process or a signal to reach a specific level or target for the first time, has attracted considerable attention [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. The classical counterpart, known as the first passage time problem, was the subject of a very large body of work; see [20, 21, 22] and references therein.
@@
For quantum walks, with monitoring, the situation is vastly different, and here we mention four such aspects.
@@
    Constructive and destructive interference. Quantum walks, for example, tight-binding models (see below), serve as a benchmark for quantum search. These walks may exhibit a quantum speedup for the hitting time, which can be exponential for specific initial states and highly symmetric graphs [2]. However, in some other cases, the quantum search can perform poorly due to destructive interference, and the latter problem can be avoided in principle using specially designed graphs [24] or with a restart strategy [25].
@@
    Path definition matters. When we obtain quantum trajectories through repeated strong measurements, the statistics of the first detection time will depend on the observation scheme with which the path is constructed. Here, we address this issue using IBM quantum computers. For that purpose, the first hitting time for tracking and for localized (on-site) measurements at the target state is defined and studied.
@@
    Topological effects. A case study is the return or recurrence problem in finite systems. The return problem addresses the time it takes for a process to return to its initial state [26, 5, 27, 10, 14]. In the classical domain, the mean recurrence time is described by Kac’s lemma [28]. This gives the mean return time in terms of the steady-state measure (see below). In a pioneering work, Grünbaum et al. [5] extended this result to the quantum domain. They showed that the mean number of measurements until the detection of the quantum walker in its initial state is quantized. This is connected to a topological effect. Mathematically, the mean return time is related to a winding number of the generating function of the first detection statistics. The mean return time is topologically protected; however, sharp transitions are found for special choices of the sampling time. These results are valid for local measurements at the target state. The authors of Ref. [29] studied a different protocol, when the trajectory is defined with a tracking protocol. In this case, the position of the quantum walker is recorded along its path until its first detection in the target state. Again, the mean return time is quantized, being larger than or equal to the one found in Ref. [5] using local monitoring.
@@
    Zeno physics. In quantum mechanics, sampling cannot be done too fast, to avoid freezing of the wave packet in its initial state [30]. Thus, the data acquisition rate is an important parameter controlling the statistics of the hitting time process.
@@ II.1 Model
When $\alpha=0$, we have a pair of distinct energy levels, and more generally when the flux $\alpha\neq 0$, we typically remove the two-fold degeneracy and the number of distinct energy levels is three.
@@ II.3 Tracking protocol
Second, we consider the tracking protocol [29], with the initial and target state |$\ket{\bf{0}}$ where we record every $\tau$ units of time the position of the walker on the graph; see Fig. 1 (right panel). In this stroboscopic protocol, measurements are made at times $\tau,2\tau,\dots,20\tau$. In a typical realization of the process, we find a result, which is given by the eigenvalues of the position operator, for example
@@
Thus, we start with state |$\ket{\psi_{\rm in}}=\ket{\bf{0}}$. Just before the first measurement at time $\tau$, the state function is |$\ket{\psi(\tau)}=U\ket{\psi_{\rm in}}$. We then measure the location of the walker on the triangle graph. Suppose that the first measurement yields $\bf{1}$, i.e., the system is projected to state |$\ket{\bf{1}}$. We then continue and evolve the system until the second measurement according to $U\ket{\bf{1}}$, and measure again.
II.4 Implementation on the Quantum Computer
@@
where $\sigma_{x}$, $\sigma_{y}$ and $\sigma_{z}$ are the Pauli matrices. The Hamiltonian $H$ defines two disconnected subspaces, the first composed of states |$\ket{00},\ket{01},\ket{10}$ and the second of |$\ket{11}$. The unitary evolution operator $U(\tau)=exp(-iH\tau)$ is decomposed into unitary gates using Cartan’s decomposition [33]. This allows us to vary $\tau$ without increasing the depth of the circuits.
@@
Now we define statistical observables of interest. They can be estimated from the sample paths and will depend on the parameters of the model, such as $\gamma\tau$, $\alpha$, the initial state and the measurement protocol. Let $F_{n}$ be the probability that the quantum walker is detected for the first time in the target state at the $n$-th measurement [8]. Thus $F_{1}$ is the probability that the target state is recorded at the first measurement event, while $F_{2}$ describes the case where the state was recorded for the first time in the second attempt. $\langle n\rangle$ is the conditional mean number of events until detection
$\langle n\rangle(N)=\frac{\sum_{n=1}^{N}nF_{n}}{\sum_{n=1}^{N}F_{n}}\ ,$ (5)
for $\sum_{n=1}^{N}F_{n}\neq 0$. Hence, $\tau\langle n\rangle$ is the mean time until detection. From now on, we will call this physical quantity mean return time for brevity. We also define the total detection probability $P_{{\rm det}}$
@@
which is the probability that a detection event was recorded. The estimation of this probability comes from the number of strings without detection. In an ideal situation with infinite resolution, these are obtained through $\lim_{N\to\infty}P_{{\rm det}}(N)=\sum_{n=1}^{\infty}F_{n}$ and if the latter is unity, that is, if the system is detected with probability one, $\lim_{N\to\infty}\langle n\rangle(N)=\sum_{n=1}^{\infty}nF_{n}$. As mentioned in the Introduction, even for the triangle model under study we can find cases where $P_{{\rm det}}<1$ due to destructive interference, as explained below.
@@ III.1 On-site protocol (theory)
The theorem of recurrence for on-site measurements reads [5]
@@
where $E_{k}$ are the eigenstates of $H$ [Eq. (3)] for the triangle model. From Eq. (3) it becomes clear why the phase factor matching diagram is a very useful tool. For example, for the choice of parameters corresponding to blue circles in Fig. 2 we expect $\langle n\rangle=1$, while for the colored lines (besides blue) $\langle n\rangle=2$, and anywhere else $\langle n\rangle=3$. This holds in general under the condition that the energy states, i.e., eigenstates of $H$, denoted |$\ket{E_{k}}$ have finite overlap with the detected state. For example, in the Zeno limit $\tau\to 0$, the three phase factors merge, and hence $\langle n\rangle$ is equal to one. This is expected for the return problem, since we start at the measured state, and hence the first measurement detects it when $\tau=0$. We provide more details in Appendix B.
@@ III.2 Tracking protocol (theory)
In general, the mean return time $\langle n\rangle$ is the number of states in the system, which is three for the triangle model under study. The result $\langle n\rangle=3$ can be viewed as a classical result. More specifically, consider a classical random walk coupled to a heat bath in the infinite temperature limit. In this case, in a steady state, all states are equally likely, namely the occupation probability is $1/3$. Then from the classical Kac theorem for the return problem $\langle n\rangle=3$ [28]. At least in the classical domain, this is easy to understand, since the probability of measuring the walker in the target state in the first measurement is $1/3$, in the second $(2/3)(1/3)$ etc. and hence $\langle n\rangle=\sum_{n=1}^{\infty}n(2/3)^{n-1}(1/3)=3$. Thus, tracking, for most of the choices of $\tau$, drives the system to a classical limit. The quantum aspect of the hitting-time process is found for special values of $\tau$ that capture the revival of the wave packet and hence the underlying periodicity of the quantum dynamics. Namely, there exist revival times where the initial wave function returns to its original state in addition to an unimportant phase. Then if we measure at that time, the quantum walker is detected with probability one in the target state in the first measurement and hence $\langle n\rangle=1$. In particular, when $\tau=0$ we have $\langle n\rangle=1$. In the diagram in Fig. 2 these special sampling times correspond to three phase factors matching, shown as blue dots and lines. Clearly, the idealized theory predicts that as we vary $\gamma\tau$ and $\alpha$, when three phase factors match, we will see a dip or a transition in $\langle n\rangle$. Thus, $\langle n\rangle$ will jump from $\langle n\rangle=3$ to $\langle n\rangle=1$ and back. In Appendix C, following Ref. [29] we provide more details of the theory by introducing the corresponding Markov or stochastic matrix.
@@ IV First hitting return times on IBM quantum computers
Theoretically, $\langle n\rangle$ exhibits pointwise discontinuous behavior as mentioned. Can we see this on a quantum computer? At first glance, this might seem hard since the width of these transitions is theoretically zero but in reality for finite $N$ the transitions are broadened.

### 10 Apr 2024 v2

# First Hitting Times on a Quantum Computer: Tracking vs. Local Monitoring, Topological Effects, and Dark States

@@ Abstract
We investigate a quantum walk on a ring represented by a directed triangle graph with complex edge weights and monitored at a constant rate until the quantum walker is detected. To this end, the first hitting time statistics is recorded using unitary dynamics interspersed stroboscopically by measurements, which is implemented on IBM quantum computers with a midcircuit readout option. Unlike classical hitting times, the statistical aspect of the problem depends on the way we construct the measured path, an effect that we quantify experimentally. First, we experimentally verify the theoretical prediction that the mean return time to a target state is quantized, with abrupt discontinuities found for specific sampling times and other control parameters, which has a well-known topological interpretation. Second, depending on the initial state, system parameters, and measurement protocol, the detection probability can be less than one or even zero, which is related to dark-state physics. Both, return-time quantization and the appearance of the dark states are related to degeneracies in the eigenvalues of the unitary time evolution operator. We conclude that, for the IBM quantum computer under study, the first hitting times of monitored quantum walks are resilient to noise. Yet, a finite number of measurements leads to broadening effects, which modify the topological quantization and chiral effects of the asymptotic theory with an infinite number of measurements. Our results point the way for the development of novel quantum walk algorithms that exploit measurement-induced effects on quantum computers.
@@ I Introduction
The option of midcircuit readout of qubits on state-of-the-art quantum computers [1] opens opportunities to test the dynamics and statistics of monitored quantum dynamics. The repeated monitoring at predetermined times yields a string of measurement outputs. These can be viewed as a stochastic trajectory varying in time which depends on many factors, including the initial state, the dynamics of the measurement-free process, or the chosen time intervals between measurements. Given a stochastic trajectory, the first quantum hitting time, defined as the time taken for a monitored process or a signal to reach a specific level or target for the first time, has attracted considerable attention [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. The classical counterpart, known as the first passage time problem, was the subject of a very large body of work; see [21, 22, 23] and references therein.
@@
For quantum walks, which are monitored by repeated projective measurements, the situation is vastly different, and here we mention four such aspects.
@@
    Constructive and destructive interference. Quantum walks, for example, tight-binding models (studied here), serve as a benchmark for quantum search [25, 26]. These walks may exhibit a quantum speedup for the hitting time, which can be exponential for specific initial states and highly symmetric graphs [2]. However, in some other cases, the quantum search can perform poorly due to destructive interference, and the latter problem can be avoided in principle using specially designed graphs [27] or with a restart strategy [28].
@@
    Path definition matters. When we obtain quantum trajectories through repeated projective measurements, the statistics of the first detection time will depend on the observation scheme with which the path is constructed. Here, we address this issue using IBM quantum computers. For that purpose, the first hitting time for tracking and for localized (on-site) measurements at the target state is defined and studied. We will also highlight when these protocols exhibit a classical, in the sense of classical random walk behaviors, and when do they exhibit quantum features.
@@
    Topological effects. A case study is the return or recurrence problem in finite systems. The return problem addresses the time it takes for a process to return to its initial state [29, 5, 6, 30, 11, 15]. In the classical domain, the mean recurrence time is described by Kac’s lemma [31]. This gives the mean return time in terms of the steady-state measure (see below). In a pioneering work, Grünbaum et al. [5, 6] extended this result to the quantum domain. They showed that the mean number of measurements until the detection of the quantum walker in its initial state is quantized. This is connected to a topological effect. Mathematically, the mean return time is related to a winding number of the generating function of the first detection statistics. One of the most striking differences between classical and quantum walks is a phase in the latter. Although the phase is usually not directly observable, its properties can have a significant impact on observable quantities. Here, the mean return time is topologically protected; however, sharp transitions are found for special choices of the sampling time. It turns out that the mean return time number is directly linked to the winding number, which implies a quantization of the latter [5, 6]. Moreover, the winding number of the return problem is equal to the dimensionality of the available Hilbert space. These results are valid for local measurements at the target state, when the information about the trajectory between measurements is unknown.
    The authors of Ref. [32] studied a different protocol, when the trajectory is defined with a tracking protocol. In this case, the position of the quantum walker is recorded along its path until its first detection in the target state. Again, the mean return time is quantized, being larger than or equal to the one found in Ref. [5, 6] using local monitoring. When monitoring obtains the information about the trajectory of the quantum walker, we destroy the phase, and hence for most sampling rates the behavior we find experimentally below is classical. However, for specific sampling rates, quantum feature are shown to emerge. In contrast, the local measurement protocol investigated here exhibits robust quantum characteristics that are independent of the sampling rate.
@@
    Zeno physics. In quantum mechanics, sampling cannot be done too fast, to avoid freezing of the wave packet in its initial state [33]. Thus, the data acquisition rate is an important parameter controlling the statistics of the hitting time process in both protocols.
@@ II.1 Model
When $\alpha=0$, we have a pair of distinct energy levels, and more generally when the flux $\alpha\neq 0$, we typically remove the two-fold degeneracy and the number of distinct energy levels is three [35].
@@ II.3 Tracking protocol
Second, we consider the tracking protocol [32], with the target state |$\ket{\bf{0}}$ where we record every $\tau$ units of time the position of the walker on the graph; see Fig. 1 (right panel). In this stroboscopic protocol, measurements are made at times $\tau,2\tau,\dots,20\tau$. In a typical realization of the process, we find a result, which is given by the eigenvalues of the position operator, for example
@@
II.4 The return problem
In the return problem the target state and the initial state are the same. This choice is also called the reccurence time problem. It will be studied with both protocols. Thus, we start with state |$\ket{\psi_{\rm in}}=\ket{\bf{0}}$. Just before the first measurement at time $\tau$, the state function is |$\ket{\psi(\tau)}=U\ket{\psi_{\rm in}}$. For the tracking protocol, we then measure the location of the walker on the triangle graph. Suppose that the first measurement yields $\bf{1}$, i.e., the system is projected to state |$\ket{\bf{1}}$. We then continue and evolve the system until the second measurement according to $U\ket{\bf{1}}$, and measure again. For the on-site protocol the new state is the projection to the other accessible states if we do not detect the quantum walker on site |$\ket{\bf 0}$. Note that when we study dark state physics, we use other initial conditions, see below.
II.5 Implementation on the Quantum Computer
@@
where $\sigma_{x}$, $\sigma_{y}$ and $\sigma_{z}$ are the Pauli matrices. The Hamiltonian $H$ defines two disconnected subspaces, the first composed of states |$\ket{00},\ket{01},\ket{10}$ and the second of |$\ket{11}$. The unitary evolution operator $U(\tau)=\exp(-iH\tau)$ is decomposed into unitary gates using Cartan’s decomposition [37]. This allows us to vary $\tau$ without increasing the depth of the circuits.
@@
II.6 Observables
Now we define statistical observables of interest. They can be estimated from the sample paths and will depend on the parameters of the model, such as $\gamma\tau$, $\alpha$, the initial state and the measurement protocol. Let $F_{n}$ be the probability that the quantum walker is detected for the first time in the target state at the $n$-th measurement [9, 32]. Thus $F_{1}$ is the probability that the target state is recorded at the first measurement event, while $F_{2}$ describes the case where the state was recorded for the first time in the second attempt. $\langle n\rangle$ is the conditional mean number of events until detection
$\langle n(N)\rangle=\frac{\sum_{n=1}^{N}nF_{n}}{\sum_{n=1}^{N}F_{n}}\ ,$ (5)
for $\sum_{n=1}^{N}F_{n}\neq 0$. Hence, $\tau\langle n(N)\rangle$ is the mean time until detection. From now on, we will call this physical quantity mean return time for brevity. We also define the total detection probability $P_{{\rm det}}$
@@ II.6 Observables
which is the probability that a detection event was recorded. The estimation of this probability comes from the number of strings without detection. In an ideal situation with infinite resolution, these are obtained through $\lim_{N\to\infty}P_{{\rm det}}(N)=\sum_{n=1}^{\infty}F_{n}$ and if the latter is unity, that is, if the system is detected with probability one, $\lim_{N\to\infty}\langle n(N)\rangle=\sum_{n=1}^{\infty}nF_{n}$. As mentioned in the Introduction, even for the triangle model under study we can find cases where $P_{{\rm det}}<1$ due to destructive interference, as explained below.
@@ III.1 On-site protocol (theory)
The theorem of recurrence for on-site measurements reads [5, 6]
@@
where $E_{k}$ are the eigenstates of $H$ [Eq. (3)] for the triangle model. From Eq. (3) it becomes clear why the phase factor matching diagram is a very useful tool. For example, for the choice of parameters corresponding to blue circles in Fig. 2 we expect $\langle n\rangle=1$, while for the colored lines (besides blue) $\langle n\rangle=2$, and anywhere else $\langle n\rangle=3$. This holds in general under the condition that the energy states, i.e., eigenstates of $H$, denoted |$\ket{E_{k}}$ have finite overlap with the detected state. For example, in the Zeno limit $\tau\to 0$, the three phase factors merge, and hence $\langle n\rangle$ is equal to one. This is expected for the return problem, since we start at the measured state, and hence the first measurement detects it when $\tau=0$. Phase factor matching conditions correspond to removal of a state from the effective Hilbert space. If phase factor match $\exp(-iE_{1}\tau)=\exp(-iE_{2}\tau)$, the state $\psi\sim\langle{\bf 0}|E_{1}\rangle|E_{2}\rangle-\langle{\bf 0}|E_{2}\rangle|E_{1}\rangle$ cannot be detected. Thus, the phase factor matching choices of $\gamma\tau$ and $\alpha$ reduce the dimension of the Hilbert space, which can be shown to lead to faster detection compared to the cases where the phase factors do not match.
We provide more details in Appendix B.
@@ III.2 Tracking protocol (theory)
Except for special model parameters, the mean return time $\langle n\rangle$ is the number of states in the system, which is three for the triangle model under study. The result $\langle n\rangle=3$ can be viewed as a classical result. More specifically, consider a classical random walk coupled to a heat bath in the infinite temperature limit. In this case, in a steady state, all states are equally likely, namely the occupation probability is $1/3$. Then from the classical Kac theorem for the return problem $\langle n\rangle=3$ [31]. At least in the classical domain, this is easy to understand, since the probability of measuring the walker in the target state in the first measurement is $1/3$, in the second $(2/3)(1/3)$ etc. and hence $\langle n\rangle=\sum_{n=1}^{\infty}n(2/3)^{n-1}(1/3)=3$. Thus, tracking, for most of the choices of $\tau$, drives the system to a classical limit. The quantum aspect of the hitting-time process is found for special values of $\tau$ that capture the revival of the wave packet and hence the underlying periodicity of the quantum dynamics. Namely, there exist revival times where the initial wave function returns to its original state in addition to an unimportant phase. Then if we measure at that time, the quantum walker is detected with probability one in the target state in the first measurement and hence $\langle n\rangle=1$. In particular, when $\tau=0$ we have $\langle n\rangle=1$. In the diagram in Fig. 2 these special sampling times correspond to three phase factors matching, shown as blue dots and blue lines. Clearly, the idealized theory predicts that as we vary $\gamma\tau$ and $\alpha$, when three phase factors match, we will see a dip or a transition in $\langle n\rangle$. Thus, $\langle n\rangle$ will jump from $\langle n\rangle=3$ to $\langle n\rangle=1$ and back. In Appendix C, following Ref. [32] we provide more details of the theory by introducing the corresponding Markov or stochastic matrix.
@@ IV First hitting return times on IBM quantum computers
Theoretically, $\langle n\rangle$ exhibits pointwise discontinuous behavior as mentioned. Can we see this on a quantum computer? At first glance, this might seem hard since the width of these transitions is theoretically zero, but in reality for finite $N$ the transitions are broadened. In the following we compare the experimental values, the classical simulation of Eq. 5 and the asymptotic theory ($N\rightarrow\infty$).
Figure 3: (a) Classical simulation of the mean return time $\langle n\rangle$ for $N=20$ for the on-site protocol as a function of $\gamma\tau$ and $\alpha$. For $\langle n\rangle=1$ three phase factors match (dark blue), for $\langle n\rangle=2$ two phase factors match (medium blue) and for $\langle n\rangle=3$ all phase factors are different (white). The paths through the parameter space labeled as (I), (II), (III) and (IV) are indicated with red dashed lines and correspond to (c), (d), (e) and (f), respectively. Compared to the phase factor matching diagram in Fig. 2 the lines are broadened due to finite $N$. (b) Quantum circuit for two qubits representing the three localized states for the on-site protocol with the initial state |$\ket{\bf{0}}=\ket{01}$, where the unitary $U=\exp{(-iH\tau)}$ and measurements are applied alternately. Only the upper qubit is measured, since we need to detect the state |$\ket{\bf{0}}$ while we do not want to receive the information that allows us to distinguish the states |$\ket{\bf{1}}$ and |$\ket{\bf{2}}$. (c) - (f) Mean return time $\langle n\rangle$ recorded on IBM Sherbrooke (light blue circles), obtained with the asymptotic theory for $N\rightarrow\infty$ (black solid lines) and simulated for $N=20$ (blue dashed lines): (c) [path (I)] $\langle n\rangle$ versus $\gamma\tau$ for $\alpha=0$. (d) [path (II)] When $\alpha=0.5$ we almost always find $\langle n\rangle=3$. As explained in the text, this is related to the removal of energy-level degeneracy when the magnetic flux is turned on. Note that some fine structure details predicted by the asymptotic theory, for $\alpha=0.5$, are washed out in the experiment, see the three nearby dips that merge into one resonance here. (e) [path (III)] Mean return time $\langle n\rangle$ as a function of $\alpha$ for $\gamma\tau=3\pi/\sqrt{3}$. $\langle n\rangle$ develops a plateau, as well as additional transitions that are absent for $N\rightarrow\infty$. (f) [path (IV)] Mean return time $\langle n\rangle$ for $\gamma\tau=2\pi/\sqrt{3}$ as a function of $\alpha$. In this example clearly the asymptotic theory is not predictive, while finite time simulations agree with experiments.
Figure 4: (a) Classical simulation of the mean return time $\langle n\rangle$ for $N=20$ for the tracking protocol as a function of $\gamma\tau$ and $\alpha$. For $\langle n\rangle=1$ three phase factors match (dark blue), and for $\langle n\rangle=3$ all phase factors are different (white). Paths through the parameter space labeled (I), (II), (III) and (IV) are indicated with red dashed lines and correspond to (c), (d), (e), and (f), respectively. Compared to the phase factor matching diagram in Fig. 2 the dark blue areas are broadened due to finite $N$. (b) Quantum circuit for two qubits representing the three localized states for the tracking protocol with the initial state |$\ket{\bf{0}}=\ket{00}$, where the unitary $U=\exp{(-iH\tau)}$ and the measurements are applied alternately. (c) - (f) Mean return time $\langle n\rangle$ recorded on IBM Sherbrooke (light blue circles), obtained with the asymptotic theory for $N\rightarrow\infty$ (black solid line) and simulated for $N=20$ (blue dashed line): (c) [path (I)] $\langle n\rangle$ versus $\gamma\tau$ for $\alpha=0$. $\langle n\rangle$ is quantized ($\langle n\rangle=3$) and its value, for almost any choice of $\gamma\tau$, differs by unity from the on-site measurement protocol, the case presented in Fig. 3 (c) when $\langle n\rangle=2$. (d) [path (II)] For finite magnetic flux $\alpha=0.5$ the transitions to $\langle n\rangle=1$ are absent except in the Zeno limit at $\gamma\tau=0$ for $N\rightarrow\infty$. This is different for $N=20$. The result for $\alpha=0.5$ shows a transition to approximately $\langle n\rangle=1$ at $\gamma\tau=3.63$. (e) [path (III)] Mean return time $\langle n\rangle$ for $\gamma\tau=3\pi/\sqrt{3}$. For $N=20$ $\langle n\rangle$ develops additional transitions that are absent for $N\rightarrow\infty$ (f) [path (IV)] Mean return time $\langle n\rangle$ for $\gamma\tau=2\pi/\sqrt{3}$ as a function of $\alpha$. We see a clear broadening effect of the resonance, while in (d) and (e) broadened resonances show up, which are not present in the asymptotic theory.